Consider two equal point charges:

(a) If two objects with a charge of 1 C each are separated in air by a distance of 1 km, what is the magnitude of the force acting between them?

(b) For comparison, the typical magnitudes of charges that we have been dealing with in the lab this quarter are much smaller (on the order of 10-7 C). If you have two equal point charges of 10-7 C, how far apart would you need to place them in order for the force between them to be the same as the force you calculated in part (a)?

(c) Describe in words what calculations in (a) and (b) tell you about the fundamental nature of electric force.
Answer #1

a)

F = k q^2 / r^2

F = 9*10^9* 1^2 / 1000^2

F = 9000 N

=======

b)

now

F = k q^2 / r^2

r^2 = 9*10^9*10^-14 / 9000

r = 10^-4 m

======

from the above two parts we can conclude that the physical magntiude of 1 C charge is too high, its really a very massive magnitude of charge that can act a force of 9kN even having a quite large Seperation distance.

======